OpenStudy (anonymous):

if x equals 8 then find the matching value for y 2x+2y=-2

OpenStudy (imstuck):

THis would be a case of solving for x and y. First replace the x with the 8 that the problem gives you. 2 times 8 is 16, so 2x + 2y = -2 becomes 16 + 2y = -2. Subtract 16 from both sides to get y alone. 2y = -2 - 16. -2 - 16 = -18; so 2y = -18. Now divide both sides by 2 to solve for y. y = -9. See that? Are you familiar with that concept in school?